# The Computational Complexity of Circuit Discovery for Inner Interpretability

### v2

@@ 4.1 Sufficient Circuit
Sufficient circuits (SCs) are sets of neurons connected end-to-end that suffice, in isolation, to reproduce some model behavior over an input domain. Therefore, they are conceptually related to the desired outcome of masking components that do not contribute to the behavior of interest (small, parameter-efficient subnetworks). SCs remain relevant as, despite valid criticisms of zero-ablation (e.g., Conmy et al., 2023), circuit discovery through pruning might be justified (Yu et al., 2024a).
Problem 1 (Bounded Locally Sufficient Circuit (BLSC))
@@
We find that many variants of SC are NP-hard (see Table 4). Counterintuitively, this intractability does not depend straightfowardly on parameters such as network depth (W[1]-hard relative to $\mathcal{P}$). Therefore, hardness is not mitigated by keeping models shallow. Given this barrier, we explore the possibility of obtaining approximate solutions but find that hard SC variants are inapproximable relative to all schemes in Section 3.3. An alternative is to consider the membership of these problems in a well-studied class whose solvers are better understood than interpretability heuristics (de Haan & Szeider, 2017). We prove that local versions of SC are NP-complete. This implies there exist efficient transformations from instances of SC to those of the satisfiability problem (SAT), opening up the possibility to borrow techniques that work reasonably well in practice for SAT (Biere et al., 2021). Interestingly, this is not possible for the global version, which we prove is complete for a class higher up the complexity hierarchy ($\Sigma^{p}_{2}$-complete). This result establishes a formal separation between local and global query complexity that partly explains ‘interpretability illusions’ (Friedman et al., 2024; Yu et al., 2024a) arising in practice due to local but not global faithfulness.
Next we explore the following alternative scenario. Given a model and a target behavior, if we knew that SCs with some desired property (e.g., minimality) are abundant, this would provide some confidence in the ability of heuristic search to stumble upon one of them. To explore this, we analyze various queries where the output is a count of the number of SCs (i.e., counting problems). We find that both local and global, bounded and unbounded variants are #P-complete and remain intractable (#W[1]-hard) when parameterized by many network features including depth (Table 3).
The hardness profile of SC over all these variants calls for exploring more substantial relaxations. We introduce the notion of quasi-minimality of SCs for this purpose and later demonstrate its usefulness beyond this particular problem. Any neuron in a minimal/minimum SC is a breaking point in the sense that removing it will break the target behavior. In quasi-minimal SCs we are guaranteed to know at least one neuron that causes this breakdown, though there may be other neurons that do not.
Problem 2 (Unbounded Quasi-minimal Locally Sufficient Circuit (UQLSC))
@@
    a circuit $\mathcal{C}$ in $\mathcal{M}$ and a neuron $v\in\mathcal{C}$ such that
    $\mathcal{C}(\mathbf{x})=\mathcal{M}(\mathbf{x})$ and $[\mathcal{C}\setminus\{v\}](\mathbf{x})\neq\mathcal{M}(\mathbf{x})$.
By introducing this relaxation, which gives up some affordances but retains others of interest, we get a feasible interpretability query. Input: is in PTIME. We describe an efficient algorithm to compute it which can be heuristically biased towards finding smaller circuits and combined with techniques that exploit weights and gradients (see The Computational Complexity of Circuit Discovery for Inner Interpretability).
@@ 4.2 Gnostic Neuron
    A set of neurons $V$ in $\mathcal{M}$ of size $|V|\geq k$ such that $\forall_{v\in V}$ it is the case that $\forall_{\mathbf{x}\in\mathcal{X}}$ computing $\mathcal{M}(\mathbf{x})$ produces activations $A^{v}_{\mathbf{x}}\geq t$ and $\forall_{\mathbf{y}\in\mathcal{Y}}:A^{v}_{\mathbf{y}}<t$,
    if it exists, otherwise $\bot$.
It is probably not a coincidence that (i) the idea and search for such neurons has emerged early and remained popular, and (ii) it is a tractable problem. Input: is in PTIME. Alternatives might require GNs to have some behavioral effect when intervened; such variants would remain tractable.
@@ 4.3 Circuit Ablation and Clamping
    a subset of neurons $\mathcal{C}$ in $\mathcal{M}$ of size $|\mathcal{C}|\leq u$, such that $[\mathcal{M}\setminus\mathcal{C}](\mathbf{x})\neq\mathcal{M}(\mathbf{x})$,
    if it exists, otherwise $\bot$.
@@
    a multi-layer perceptron $\mathcal{M}$, an input vector $\mathbf{x}$, a value $r$, and an integer $u$ with $1<u\leq|\mathcal{M}|$.
@@
Despite these more modest criteria, we find that both the local and global variants of CA and CC are NP-hard, fixed-parameter intractable W[1]-hard relative to various parameters, and inapproximable in all 5 senses studied. However, we prove these problems are in NP-complete, which opens up practical options not available for other problems we study (see remarks in Section 4.1).
@@ 4.4 Circuit Patching
A critique of zero-ablation is the arbitrariness of the value, leading to alternatives such as mean-ablation (e.g., Wang et al., 2022). A related contrast is studying circuits in isolation versus embedded in surrounding subnetworks. Activation patching (Ghandeharioun et al., 2024; Zhang & Nanda, 2023) and path patching (Goldowsky-Dill et al., 2023) try to pinpoint which activations play an in-context role in model behavior over a domain, which inspires the circuit patching (CP) problem.
@@
    a multi-layer perceptron $\mathcal{M}$, an integer $k$, an input vector $\mathbf{y}$, and a set $\mathcal{X}$ of input vectors.
@@
In this way we obtain a tractable query (PTIME) for quasi-minimal patching, sidestepping barriers while retaining some useful affordances (see Table 1). We present an algorithm to compute Input: efficiently that can be combined with strategies exploiting weights and gradients (see The Computational Complexity of Circuit Discovery for Inner Interpretability).
@@
Classical &
parameterized problems [^2] Problem versions are for bounded size circuits unless otherwise stated. Each cell contains the complexity of the problem variant in terms of classical and FP (in)tractability, membership in complexity classes, and various approximation schemes. ‘?’ marks potentially fruitful open problems. ‘N/A’ stands for not applicable.
$\mathcal{P}_{\mathcal{M}}=\{\hat{L},\hat{U}_{I},\hat{U}_{O},\hat{W},\hat{B}\}\newline \mathcal{P}_{\mathcal{C}}=\{\hat{l},\hat{l}_{w},\hat{u},\hat{u}_{I},\hat{u}_{O},\hat{w},\hat{b}\}$
@@
Sufficient Circuit (SC) NP-complete {$c$, PTAS, 3PA}-inapproximable $\Sigma^{p}_{2}$-complete {$c$, PTAS, 3PA}-inapproximable
$\mathcal{P}$-SC
$(\mathcal{P}=\mathcal{P}_{\mathcal{M}}\cup\mathcal{P}_{\mathcal{C}})$
W[1]-hard {$c$, PTAS, 3PA}-inapproximable W[1]-hard {$c$, PTAS, 3PA}-inapproximable
@@
Unbounded Minimal SC ? N A $\in\Sigma^{p}_{2}$ NP-hard N A
@@
Circuit Ablation (CA) NP-complete {$c$, PTAS, 3PA}-inapproximable $\in\Sigma^{p}_{2}$ NP-hard {$c$, PTAS, 3PA}-inapproximable
$\{\hat{L},\hat{U}_{I},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-CA W[1]-hard {$c$, PTAS, 3PA}-inapproximable W[1]-hard {$c$, PTAS, 3PA}-inapproximable
Circuit Clamping (CC) NP-complete {$c$, PTAS, 3PA}-inapproximable $\in\Sigma^{p}_{2}$ NP-hard {$c$, PTAS, 3PA}-inapproximable
$\{\hat{L},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-CC W[1]-hard {$c$, PTAS, 3PA}-inapproximable W[1]-hard {$c$, PTAS, 3PA}-inapproximable
Circuit Patching (CP) NP-complete {$c$, PTAS, 3PA}-inapproximable $\in\Sigma^{p}_{2}$ NP-hard {$c$, PTAS, 3PA}-inapproximable
$\{\hat{L},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-CP W[2]-hard {$c$, PTAS, 3PA}-inapproximable W[2]-hard {$c$, PTAS, 3PA}-inapproximable
@@
Necessary Circuit (NC) $\in\Sigma^{p}_{2}$ NP-hard {$c$, PTAS, 3PA}-inapproximable $\in\Sigma^{p}_{2}$ NP-hard {$c$, PTAS, 3PA}-inapproximable
$\{\hat{L},\hat{U}_{I},\hat{U}_{O},\hat{W},\hat{u}\}$-NC W[1]-hard {$c$, PTAS, 3PA}-inapproximable W[1]-hard {$c$, PTAS, 3PA}-inapproximable
Count NC ? ? ? ?
@@
Sufficient Reasons (SR) $\in\Sigma^{p}_{2}$ NP-hard {3PA}-inapproximable N A
$\{\hat{L},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-SR W[1]-hard {3PA}-inapproximable
@@ 4.5 Necessary Circuit
The criterion of necessity is a stringent one, and consequently necessary circuits (NCs) carry powerful affordances (see Table 1). Since neurons in NCs collectively interact with all possible sufficient circuits for a target behavior, they are candidates to describe key task subcomputations and intervening on them is guaranteed to have effects even in the presence of high redundance.
Problem 8 (Bounded Globally Necessary Circuit (BGNC))
@@
Unfortunately both local and global versions of NC are NP-hard (in $\Sigma^{p}_{2}$; Table 4), remain intractable even when keeping parameters such as network depth, number of input and output neurons, and others small (Table 3), and does not admit any of the available approximation schemes (Section 3.3). Tractable versions of NC are unlikely unless substantial restrictions or relaxations are introduced.
@@ 4.6 Circuit Robustness
A behavior of interest might be over-determined or resilient in the sense that many circuits in the model implement it and one can take over when the other breaks down. This is related to the notion of redundancy used in neuroscience (e.g., Nanda et al., 2023). Intuitively, when a model implements a task in this way, the behavior should be more robust to a number of perturbations. The possibility of verifying it experimentally motivates the circuit robustness (CR) problem.
@@
    <YES> if for each subset $H^{\prime}\subseteq H$, with $|H^{\prime}|\leq k$, $\mathcal{M}(\mathbf{x})=[\mathcal{M}\setminus H^{\prime}](\mathbf{x})$,
    otherwise <NO>.
@@ 5 Implications, limitations, and future directions
We presented a framework based on parameterized complexity to accompany experiments on inner interpretability with theoretical explorations of viable algorithms. With this grasp of circuit query complexity, we can understand the challenges of scalability and the mixed outcomes of experiments with heuristics for circuit discovery. We can explain ‘interpretability illusions’ (Friedman et al., 2024) due to lack of faithfulness, minimality (e.g., Shi et al., 2024; Yu et al., 2024a) and other affordances (Wang & Veitch, 2024; Hase et al., 2023), in terms of the kinds of circuits that our current heuristics are well-equipped to discover. For instance, consider the algorithm for automated circuit discovery proposed by Conmy et al. (2023), which eliminates one network component at a time if the consequence on behavior is reasonably small. Since this algorithm runs in polynomial time, it is not likely to solve the problems proven hard here, such as Minimal Sufficient Circuit. However, one reason we observe interesting results in some cases is because it is well-equipped to solve Quasi-Minimal Circuit problems. As our conceptual and formal analyses show, quasi-minimal circuits can mimic various desirable aspects of sufficient circuits (Table 1), and the former can be found tractably (results for Input: and Input:). Indeed, from our proof of tractability of QMSC (see The Computational Complexity of Circuit Discovery for Inner Interpretability) we get a hint on how to improve the running time of the algorithm in Conmy et al. (2023) while retaining guarantees of quasi-minimality; namely, by using a variant of binary search to cut the number of forward passes from $n$ to $\log_{2}(n)$. At the same time, understanding these properties of circuit discovery heuristics helps us explain observed discrepancies: why we often see (1) lack of faithfulness (i.e., global coverage is out of reach for QMC algorithms), (2) non-minimality (i.e., QM circuits can have many non-breaking points), and (3) large variability in performance across tasks and analysis parameters (e.g., Shi et al., 2024; Conmy et al., 2023).
Our results for inner interpretability complement those of explainability (e.g., Barceló et al., 2020; Bassan et al., 2024; Wäldchen et al., 2021). These two complexity aspects can be studied together for different architectures to assess their intrinsic interpretability. As in explainability, our findings suggest interpretability methods can be sought relative to additional query restrictions or relaxations. We have explored many and some yield tractability while retaining affordances of practical interest.
@@
One possibly fruitful avenue we have not explored is to conduct an empirical and formal characterization of learned weights in search of structure that could potentially distinguish conditions of (in)tractability. Another avenue is to design queries that partially rely on mid-level abstractions (Vilas et al., 2024a) to bridge the gap between circuits and human-intelligible algorithms (e.g., key-value mechanisms; Geva et al., 2022; Vilas et al., 2024b).

### v3

@@ 4.1 Sufficient Circuit
Sufficient circuits (SCs) are sets of neurons connected end-to-end that suffice, in isolation, to reproduce some model behavior over an input domain (see faithfulness; Wang et al., 2022; Yu et al., 2024a). They are conceptually related to the desired outcome of zero-ablating components that do not contribute to the behavior of interest (small, parameter-efficient subnetworks). Zero-ablation as a method (e.g., to find sufficient circuits) has been criticized on the grounds that the patched value (zero) is somewhat arbitrary and therefore can mischaracterize the functioning of the neuron/circuit when operating in the context of the rest of the network during inference. This gives rise to alternative methods such as activation patching with activation means or specific input activations, which we study later. SCs remain relevant as, despite valid criticisms of zero-ablation (e.g., Conmy et al., 2023), circuit discovery through pruning might be justified at least in some cases (Yu et al., 2024a).
Problem 1 (Bounded Local Sufficient Circuit (BLSC))
@@
We find that many variants of SC are NP-hard (see Table 4). Counterintuitively, this intractability does not depend straightforwardly on parameters such as network depth (W[1]-hard relative to $\mathcal{P}$). Therefore, hardness is not mitigated by keeping models shallow. Given this barrier, we explore the possibility of obtaining approximate solutions but find that hard SC variants are inapproximable relative to all schemes in Table 3. An alternative is to consider the membership of these problems in a well-studied class whose solvers are better understood than interpretability heuristics (de Haan & Szeider, 2017). We prove that local versions of SC are NP-complete. This implies there exist efficient transformations from instances of SC to those of the satisfiability problem (SAT; Biere et al., 2021), opening up the possibility to borrow techniques that work reasonably well in practice for SAT that might be suitable for some versions of neural network problems. Interestingly, this is not possible for the global version, which we prove is complete for a class higher up the complexity hierarchy ($\Sigma^{p}_{2}$-complete). This result establishes a formal separation between local and global query complexity that partly explains ‘interpretability illusions’ (Friedman et al., 2024; Yu et al., 2024a), which we conjecture holds for other queries we investigate later. These illusions come about when an interpretability abstraction (e.g., a circuit) seems empirically faithful to its target (e.g., model behavior) by some criterion (e.g., ‘local’ tests on a dataset), but actually lacks faithfulness in the way it generalizes to other criteria (e.g., tests of its ‘global’ behavior outside the original distribution).
Next we explore whether we could diagnose if SCs with some desired property (e.g., minimality) are abundant, which would be informative of the ability of heuristic search to stumble upon one of them. We analyze various queries where the output is a count of SCs (i.e., counting problems). We find that both local and global, bounded and unbounded variants are #P-complete and remain intractable (#W[1]-hard) when parameterized by many network features including depth (Table 3).
The hardness profile of SC over all these variants calls for exploring more substantial relaxations. We introduce the notion of quasi-minimality for this purpose (similar to Ramaswamy, 2019) and later demonstrate its usefulness beyond this particular problem. Any neuron in a minimal/minimum SC is a breaking point in the sense that removing it will break the target behavior. In quasi-minimal SCs we are merely guaranteed to know at least one neuron that causes this breakdown. By introducing this relaxation, which gives up some affordances but retains others of interest, we get a feasible interpretability query.
Problem 2 (Unbounded Quasi-minimal Local Sufficient Circuit (UQLSC))
@@
    A circuit $\mathcal{C}$ in $\mathcal{M}$ and a neuron $v\in\mathcal{C}$ s.t. $\mathcal{C}(\mathbf{x})=\mathcal{M}(\mathbf{x})$ and $[\mathcal{C}\setminus\{v\}](\mathbf{x})\neq\mathcal{M}(\mathbf{x})$.
Input: is in PTIME. We describe an efficient algorithm to compute it which can be heuristically biased towards finding smaller circuits and combined with techniques that exploit weights and gradients (see Appendix: Definitions, Theorems and Proofs).
@@ 4.2 Gnostic Neuron
    A set of neurons $V$ in $\mathcal{M}$ of size $|V|\geq k$ such that $\forall_{v\in V}$ it is the case that $\forall_{\mathbf{x}\in\mathcal{X}}$, $\mathcal{M}(\mathbf{x})$ produces activations $A^{v}_{\mathbf{x}}\geq t$ and $\forall_{\mathbf{y}\in\mathcal{Y}}:A^{v}_{\mathbf{y}}<t$, if it exists, else $\bot$.
Input: is in PTIME. Alternatives might require GNs to have some behavioral effect when intervened; such variants would remain tractable.
@@ 4.3 Circuit Ablation and Clamping
    A neuron set $\mathcal{C}$ in $\mathcal{M}$ of size $|\mathcal{C}|\leq u$, s.t. $[\mathcal{M}\setminus\mathcal{C}](\mathbf{x})\neq\mathcal{M}(\mathbf{x})$, if it exists, else $\bot$.
@@
    A multi-layer perceptron $\mathcal{M}$, vector $\mathbf{x}$, value $r$, and an integer $u$ s.t. $1<u\leq|\mathcal{M}|$.
@@
Despite these more modest criteria, we find that both the local and global variants of CA and CC are NP-hard, fixed-parameter intractable W[1]-hard relative to various parameters, and inapproximable in all 5 senses studied. However, we prove these problems are NP-complete, which opens up practical options not available for other problems we study (see remarks in Section 4.1).
@@ 4.4 Circuit Patching
A critique of zero-ablation is the arbitrariness of the value, leading to alternatives such as mean-ablation (e.g., Wang et al., 2022). This contrasts studying circuits in isolation versus embedded in surrounding subnetworks. Activation patching (Ghandeharioun et al., 2024; Zhang & Nanda, 2023; Hanna et al., 2024) and path patching (Goldowsky-Dill et al., 2023) try to pinpoint which activations play an in-context role in model behavior, which inspires the circuit patching (CP) problem.
@@
    A multi-layer perceptron $\mathcal{M}$, an integer $k$, an input vector $\mathbf{y}$, and a vector set $\mathcal{X}$.
@@
In this way we obtain a tractable query (PTIME) for quasi-minimal patching, sidestepping barriers while retaining some useful affordances (see Table 1). We present an algorithm to compute Input: efficiently that can be combined with strategies exploiting weights and gradients (see Appendix: Definitions, Theorems and Proofs).
@@
[…]
$\mathcal{P}=\mathcal{P}_{\mathcal{M}}\cup\mathcal{P}_{\mathcal{C}}$
$\mathcal{P}_{\mathcal{M}}=\{\hat{L},\hat{U}_{I},\hat{U}_{O},\hat{W},\hat{B}\}$
$\mathcal{P}_{\mathcal{C}}=\{\hat{l},\hat{l}_{w},\hat{u},\hat{u}_{I},\hat{u}_{O},\hat{w},\hat{b}\}$
@@
Sufficient Circuit (SC) NP-complete $\mathcal{A}$-inapprox. $\Sigma^{p}_{2}$-complete $\mathcal{A}$-inapprox.
$\mathcal{P}$-SC W[1]-hard $\mathcal{A}$-inapprox. W[1]-hard $\mathcal{A}$-inapprox.
@@
Unbounded Minimal SC ? N A ? N A
@@
Circuit Ablation (CA) NP-complete $\mathcal{A}$-inapprox. $\in\Sigma^{p}_{2}$   NP-hard $\mathcal{A}$-inapprox.
$\{\hat{L},\hat{U}_{I},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-CA W[1]-hard $\mathcal{A}$-inapprox. W[1]-hard $\mathcal{A}$-inapprox.
Circuit Clamping (CC) NP-complete $\mathcal{A}$-inapprox. $\in\Sigma^{p}_{2}$   NP-hard $\mathcal{A}$-inapprox.
$\{\hat{L},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-CC W[1]-hard $\mathcal{A}$-inapprox. W[1]-hard $\mathcal{A}$-inapprox.
Circuit Patching (CP) NP-complete $\mathcal{A}$-inapprox. $\in\Sigma^{p}_{2}$   NP-hard $\mathcal{A}$-inapprox.
$\{\hat{L},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-CP W[2]-hard $\mathcal{A}$-inapprox. W[2]-hard $\mathcal{A}$-inapprox.
@@
Necessary Circuit (NC) $\in\Sigma^{p}_{2}$  NP-hard $\mathcal{A}$-inapprox. $\in\Sigma^{p}_{2}$   NP-hard $\mathcal{A}$-inapprox.
$\{\hat{L},\hat{U}_{I},\hat{U}_{O},\hat{W},\hat{u}\}$-NC W[1]-hard $\mathcal{A}$-inapprox. W[1]-hard $\mathcal{A}$-inapprox.
@@
Sufficient Reasons (SR) $\in\Sigma^{p}_{2}$   NP-hard 3PA-inapprox. N A
$\{\hat{L},\hat{U}_{O},\hat{W},\hat{B},\hat{u}\}$-SR W[1]-hard 3PA-inapprox.
@@ 4.5 Necessary Circuit
The criterion of necessity is a stringent one, and consequently necessary circuits (NCs) carry powerful affordances (see Table 1). Since neurons in NCs collectively interact with all possible sufficient circuits for a target behavior, they are candidates to describe key task subcomputations and intervening on them is guaranteed to have effects even in the presence of high redundance. This relates to the notion of circuit overlap and therefore to efforts in identifying circuits shared by various tasks (e.g., Merullo et al., 2024), and the link between overlap and faithfulness (e.g., Hanna et al., 2024).
Problem 8 (Bounded Global Necessary Circuit (BGNC))
@@
Unfortunately both local and global versions of NC are NP-hard (in $\Sigma^{p}_{2}$; Table 4), remain intractable even when keeping parameters such as network depth, number of input and output neurons, and others small (Table 3), and does not admit any of the available approximation schemes (Table 3). Tractable versions of NC are unlikely unless substantial restrictions or relaxations are introduced.
@@ 4.6 Circuit Robustness
A behavior of interest might be over-determined or resilient in the sense that many circuits in the model implement it and one can take over when the other breaks down. This is related to the notion of redundancy used in neuroscience (e.g., Nanda et al., 2023). Intuitively, when a model implements a task in this way, the behavior should be more robust to a number of perturbations. The possibility of verifying this property experimentally motivates the circuit robustness (CR) problem, and a related interpretability effort is diagnosing nodes that are excluded by circuit discovery procedures but still have an impact on behavior (false negatives; Kramár et al., 2024).
@@
    <YES> if for each $H^{\prime}\subseteq H$, with $|H^{\prime}|\leq k$, $\mathcal{M}(\mathbf{x})=[\mathcal{M}\setminus H^{\prime}](\mathbf{x})$, else <NO>.
@@ 5 Implications, limitations, and future directions
We presented a framework based on parameterized complexity to accompany experiments on inner interpretability with theoretical explorations of viable algorithms. With this grasp of circuit query complexity, we can understand the challenges of scalability and the mixed outcomes of experiments with heuristics for circuit discovery. There is ample complexity-theoretic evidence that there is a limit (often underestimated) to how good the performance of heuristics on intractable problems can be (Hemaspaandra & Williams, 2012). We can explain ‘interpretability illusions’ (Friedman et al., 2024) due to lack of faithfulness, minimality (e.g., Shi et al., 2024; Yu et al., 2024a) and other affordances (Wang & Veitch, 2024; Hase et al., 2023), in terms of the kinds of circuits that our current heuristics are well-equipped to discover. For instance, consider the algorithm for automated circuit discovery proposed by Conmy et al. (2023), which eliminates one network component at a time if the consequence on behavior is reasonably small. Since this algorithm runs in polynomial time, it is not likely to solve the problems proven hard here, such as Minimal Sufficient Circuit. However, one reason we observe interesting results in some cases is because it is well-equipped to solve Quasi-Minimal Circuit problems. As our conceptual and formal analyses show, quasi-minimal circuits can mimic various desirable aspects of sufficient circuits (Table 1), and the former can be found tractably (results for Input: and Input:). At the same time, understanding these properties of circuit discovery heuristics helps us explain observed discrepancies: why we often see (1) lack of faithfulness (i.e., global coverage is out of reach for QMC algorithms), (2) non-minimality (i.e., QM circuits can have many non-breaking points), and (3) large variability in performance across tasks and analysis parameters (e.g., Shi et al., 2024; Conmy et al., 2023).
Although we find that many queries of interest are intractable in the general case (and empirical results are in line with this characterization), this should not paralyze efforts to interpret neural network models. As our exploration of the current complexity landscape shows, reasonable relaxations, restrictions and problem variants can yield tractable queries for circuits with useful properties. Consider a few out of many possible avenues to continue these explorations.
(i) Study query parameters. Faced with an intractable query, we can investigate which parameters of the problem (e.g., network or circuit aspects) might be responsible for its core hardness. If these problematic parameters can be kept small in real-world applications, this yields a fixed-parameter tractable query. We have explored some, but more are possible as any aspect of the problem can be parameterized. A close dialogue between theorists and experimentalists is important for this, as empirical regularities suggest which parameters might be fruitful to explore theoretically, and experiments test whether theoretically conjectured parameters can be kept small in practice.
(ii) Generate novel queries. Our formalization of quasi-minimal circuit problems illustrates the search for viable algorithmic options with examples of tractable problems for inner interpretability. When the use case is well defined, efficient queries that return circuits with useful affordances for applications can be designed. Alternative circuits might also mimic the affordances for prediction/control of ideal circuits while avoiding intractability.
(iii) Explore network output as axis of approximation. Some of our constructions use binary input/output (following previous work; e.g., Bassan et al., 2024; Barceló et al., 2020). Although continuous output does not necessarily matter complexity-wise (see Appendix: Definitions, Theorems and Proofs for [counter]examples), this is an interesting direction for future work, as it opens the door to studying the network output as an axis of approximation, which in turn might be a useful relaxation.
@@
(iv) Design more abstract queries. A different path is to design queries that partially rely on mid-level abstractions (Vilas et al., 2024a) to bridge the gap between circuits and human-intelligible algorithms (e.g., key-value mechanisms; Geva et al., 2022; Vilas et al., 2024b).
(v) Characterize actual network structure. It is in principle possible that some real-world, trained neural networks possess internal structure that is benevolent to general (ideal) circuit queries (e.g., redundancy; see Appendix: Definitions, Theorems and Proofs). In such optimistic scenarios, general-purpose heuristics might work well. The empirical evidence available to date, however, speaks against this. In any case, it will always be important to characterize any such structure to use it explicitly to design algorithms with useful guarantees.
(vi) Compare resource demands of interpretability/explainability across architectures. Our results for inner interpretability complement those of explainability (e.g., Barceló et al., 2020; Bassan et al., 2024; Wäldchen et al., 2021). These aspects can be studied together for different architectures to assess their intrinsic interpretability. To some extent our results already transfer to some cases of interest. Since the transformer architecture contains MLPs, the complexity status of our circuit queries bears on neuron-level circuit discovery efforts in transformers (e.g., Large Language/Vision/Audio models).